OpenStudy (anonymous):

Sam can not read a part of this homework problem. -7y + 5(3 + __ y) Sam knows the expression simplifies to 3y + 15. What number belongs on the blank line in the homework problem?

OpenStudy (anonymous):

A. -4 B. 2 C. 4 D. 10

OpenStudy (anonymous):

Thanks for the answers

OpenStudy (anonymous):

What does it equal to

OpenStudy (anonymous):

I don't know. -.-

jimthompson5910 (jim_thompson5910):

-7y + 5(3 + __ y) -7y + 5(3 + n*y) -7y + 15 + 5ny (-7y + 5ny) + 15 (-7+5n)y + 15 This expression is equal to 3y + 15, so (-7+5n)y + 15 = 3y + 15 (-7+5n)y = 3y -7 + 5n = 3 keep going to solve for n